### v2

# On Bits and Bandits: Quantifying the Regret-Information Trade-off

Abstract
In interactive decision-making tasks, information can be acquired by direct interactions, through receiving indirect feedback, and from external knowledgeable sources. We examine the trade-off between the information an agent accumulates and the regret it suffers. We show that information from external sources, measured in bits, can be traded off for regret, measured in reward. We invoke information-theoretic methods for obtaining regret lower bounds, that also allow us to easily re-derive several known lower bounds. We then generalize a variety of interactive decision-making tasks with external information to a new setting. Using this setting, we introduce the first Bayesian regret lower bounds that depend on the information an agent accumulates. These lower bounds also prove the near-optimality of Thompson sampling for Bayesian problems. Finally, we demonstrate the utility of these bounds in improving the performance of a question-answering task with large language models, allowing us to obtain valuable insights.
1 Introduction
In interactive decision-making tasks, an agent interacts with an environment by sequentially choosing decisions from a decision space. Subsequently, the agent receives feedback that usually includes a reward and, optionally, other types of information. For example, in multi-armed bandits (MAB) [Lattimore and Szepesvári, 2020], the feedback includes only the reward. In partial monitoring, however, it only includes a signal, which the agent then needs to use to indirectly infer the reward [Cesa-Bianchi et al., 2006]. The goal of the agent in such tasks is to minimize the gap between the accumulated reward and the reward of the best decision in hindsight, also known as regret.
In these tasks, agents can accumulate information from multiple sources, that can be categorized into two main types. One is information acquired through direct interactions and receiving feedback from the task. For example, in reinforcement learning such information includes visited states, accumulated rewards, and performed actions. The other type is information provided by external sources. Such information includes human advice [Najar and Chetouani, 2021], text description generated by large language models [Du et al., 2023, Hu and Sadigh, 2023], and more. Figure 1 describes a general decision-making task with external information. Resource allocation is an example of a task that combines both types of information. In this task, our goal is to route a process from a mobile device to a server while minimizing energy consumption and execution time [Zuo and Joe-Wong, 2021]. In this task, we can accumulate information directly by routing processes and receiving feedback. However, we also have relevant information that can be obtained from other sources, such as server specifications, load information, and more.
When we receive information from multiple sources, we can measure the total information in bits. However, we assess the algorithm’s performance using regret. Thus, a question arises:
What is the exact relationship between information bits and regret?
Figure 1: Schematic of general interactive decision-making with contextual information task. Every round, a source of knowledge provides our agent with information. The agent then makes a decision, that causes the task to generate an observation and a reward. The observation is revealed to the agent, who updates his next decision according to the past observations and the information he received.
1.1 Contributions
New method for obtaining regret lower bounds.
We introduce a novel information-theoretic method to obtain regret lower bounds for a wide variety of interactive decision-making problems. This method uses Fano’s inequality [Cover, 1999] and the construction method introduced by Yang and Barron [1999] to lower bound the worst-case Bayesian regret. We then use this method to derive lower bounds for multi-armed bandits, contextual bandits, and reinforcement learning tasks. This method allows us to lower-bound both the Bayesian and frequentist regret.
Information-theoretic prior-dependent lower bounds
can be obtained using this method. We consider an interactive decision-making framework that describes general online tasks with contextual information in a Bayesian setting. We examine the regret of policies with a mutual information constraint and obtain Bayesian regret lower bounds that depend on the entropy of our prior. These bounds also show that both Thompson sampling [Russo and Van Roy, 2016] and information-directed sampling (IDS) [Russo and Van Roy, 2014] are nearly optimal in the Bayesian sense for multi-armed and linear bandits.
Regret-information trade-off.
We use the lower bounds to quantify the relationship between information, measured in bits, and regret. We show that the presented lower bound, combined with existing upper bounds, quantifies the relationship between external information and regret. Furthermore, these bounds quantify the relationship between information accumulation and regret in online tasks. Finally, we utilize the trade-off in a question-answering task with large language models, demonstrating how we can easily utilize this trade-off in online settings.
2 Setting and Preliminaries
We introduce a new setting called interactive decision-making in Bayesian environments, with contextual information. This setting is a Bayesian adaptation of the frequentist interactive decision-making setting introduced by Foster et al. [2021]. We denote the set of Borel probability measures over a locally compact space by $\Delta(\cdot)$. Let $\Pi$ be a compact decision space and $\mathcal{M}$ be a compact class of models, where each model $M\in\mathcal{M}$ is a mapping from the decision space to an observation space $\mathcal{O}$, i.e., $M:\Pi\to\mathcal{O}$. Furthermore, every model $M$ has a reward function, $R_{M}:\Pi\to\Delta([0,1])$ that maps every decision to a reward distribution over $[0,1]$. $\mathcal{C}$ is another compact space called the context space. A task in this setting is defined by the decision space $\Pi$, the model class $\mathcal{M}$, the context space $\mathcal{C}$, and the prior $P\in\Delta(\mathcal{M}\otimes\mathcal{C})$. We denote by $\mu_{M}(\pi)=\mathbb{E}_{x\sim R_{M}(\pi)}[x]$ the mean reward of decision $\pi$ under model $M$. With a slight abuse of notation, we also denote by $\mu_{M}(\phi)=\mathbb{E}_{\pi\sim\phi}[\mu_{M}(\pi)]$, the mean reward of a policy $\phi\in\Delta(\Pi)$ under model $M$.
In a $T$-round game, a random model and context $(M_{t},C_{t})$ are sampled every round, independently from previous rounds, according to the prior probability $P$. Before each round starts, the context $C_{t}$ is revealed to the agent, who determines his policy $p_{t}\in\Delta(\Pi)$ according to the context and the observed history. Then, a decision $\pi_{t}\sim p_{t}$ is sampled and an observation $O_{t}\sim M_{t}(\pi_{t})$ is revealed to the agent. A reward, $r_{t}\sim R_{M_{t}}(\pi_{t})$ is also generated. We denote the history up to time $t$ as $\mathcal{H}_{t}=\{C_{i},\pi_{i},O_{i}\}_{i=1}^{t}$. We measure the performance with respect to the best context-dependent decision in hindsight. This performance metric is called the Bayesian regret and is defined by
$\mathcal{BR}_{P}(T;\{\pi_{t}\}_{t=1}^{T})=\sup_{\pi^{*}:\mathcal{C}\to\Pi}\mathbb{E}\left[\sum_{t=1}^{T}\left(R_{M_{t}}(\pi^{*}({c_{t}}))-R_{M_{t}}(\pi_{t})\right)\right]$
where the expectation is taken with respect to the randomness in decisions, observations, and rewards ($(M_{t},C_{t})\sim P$, $\pi_{t}\sim p_{t}$ and $r_{t}\sim R_{M_{t}}(\pi_{t})$). We also note that by fixing a single context, $\left|\mathcal{C}\right|=1$, the setting reduces to regular Bayesian interactive decision-making. When obvious from the context, $P$ and $\{\pi_{t}\}_{t=1}^{T}$ are omitted. We now show how our setting covers common decision-making tasks with contextual information.
Example 2.1 (Contextual multi-armed bandit with Bernoulli rewards).
We demonstrate how our general framework reduces to the contextual MAB problem, where the rewards for each arm are Bernoulli random variables. Assume a finite set of decisions, which are called arms in this setting, $\Pi=\{1,2,\dots,A\}$. Furthermore, assume that the context space is finite as well, $\mathcal{C}=\{1,2,\dots,C\}$. The model class is then given by =$\mathcal{M}=\{f|f:\Pi\times\mathcal{C}\to[0,1]\}$, so every model maps $\left(\pi,C\right)$ to $\mathrm{Bern}(f(\pi,C))$. At each round $t$, a model and context are sampled $(M_{t},C_{t})\sim P$, and the context is revealed. A decision $\pi_{t}$ is selected and the observation is the incurred reward, $r_{t}=O_{t}\sim\mathrm{Bern}(f_{t}(\pi_{t},C_{t}))$.
Example 2.2 (Tabular reinforcement learning with a finite horizon).
We also demonstrate how our setting applies to a finite horizon MDP. We consider a tabular MDP with a finite horizon, which is defined according to the tuple $(\mathcal{S},\mathcal{A},\mathbb{P},\mathbb{R},H)$. We also assume the episodic and stationary setting, where we update the policy only at the beginning of every episode. Additionally, we assume that the reward of every episode is bounded in $[0,1]$. In this setting, the decision space is all of the deterministic policies, so $\Pi=\{\pi:\mathcal{S}\times\mathcal{H}\to\mathcal{A}\}$. The model class $\mathcal{M}$ is defined by the set of all MDPs with the same state space, action set, and horizon. At every time step, the agent determines a policy and plays it over the entire horizon. The observation is the trajectories, which include the visited states, rewards, and actions performed.
Covering and packing
Assume a normed space $A$ with metric $d$ induced by the norm, and let $\epsilon>0$. A set $F\subseteq A$ is called a covering set of $A$ if for all $a\in A$, there exists $f\in F$ such that $d(a,f)\leq\epsilon$. The $\epsilon$-covering number of $A$, denoted by $\mathcal{N}(d,A,\epsilon)$, is the smallest cardinality possible for an $\epsilon$-covering set. A set $G\subseteq A$ is called a packing set of $A$ if $\forall g_{1},g_{2}\in G$, $d(g_{1},g_{2})\geq\epsilon$. The $\epsilon$-packing number of $A$ is the largest possible cardinality for an $\epsilon$-packing set. We define an $\epsilon$-local set of A as $B(b,\epsilon)=\{a\in A:d(b,a)\leq\epsilon\}$, for every $b\in A$ and $\epsilon>0$. The local packing number, denoted by $\mathcal{M}^{\mathrm{local}}\left(d,A,\epsilon\right)$, is the largest $(\epsilon/2)$-packing set of any set $B(b,\epsilon)$, i.e.,
$\mathcal{M}^{\mathrm{local}}\left(d,A,\epsilon\right)=\max\left\{M:\exists b\text{ such that the }(\epsilon/2)\text{-packing number of }B(b,\epsilon)\text{ is }M\right\}.$
A set $E\subseteq A$ will be called an $\epsilon$-local packing set if $\exists b$ such that $E$ is an $(\epsilon/2)$-packing set of $B(b,\epsilon)$ and $\left|E\right|=\mathcal{M}^{\mathrm{local}}\left(d,A,\epsilon\right)$. When $d(x,y)=\left\|x-y\right\|_{p}$ we denote $\mathcal{M}_{p}(A,\epsilon),\mathcal{N}_{p}(A,\epsilon)$ and $\mathcal{M}^{\mathrm{local}}_{p}\left(A,\epsilon\right)$ for the $\epsilon$-packing, $\epsilon$-covering and $\epsilon$-local packing number, respectively.
Additional notations
We denote by $H(X)=\mathbb{E}\left[-\log Q(X)\right]$ the Shannon entropy of a random variable $X\sim Q$, and with a slight abuse of notation also denote $H(Q)=H(X)$. Given two jointly distributed random variables, $(X,Y)\sim Q$, we denote their marginal distributions by $Q(x)$ and $Q(y)$. We also define the mutual information of two jointly distributed random variables $(X,Y)\sim Q$ as $I(X;Y)=\mathbb{E}\left[\log\left(\frac{Q(X,Y)}{Q(X)Q(Y)}\right)\right]$. The KL-divergence is defined as $D\left(P,Q\right)=\mathbb{E}_{X\sim P}\log\frac{P(X)}{Q(X)}$ Unless stated otherwise, logarithms are assumed to be in base-2, denoted by $\log$. We also denote $x\gtrsim y$ if there is some global constant $c>0$ such that $x\geq cy$.
Table 1: Lower bounds for MAB and RL problems, recovered with Theorem 3.5 and Proposition 4.3.
Problem Retrieved lower bound Optimal?
MAB $\Omega(\sqrt{KT})$ Yes
Tabular, episodic and stationary MDP $\Omega(\sqrt{HSKT})$ Yes [Zhang et al., 2021]
Linear bandits $\Omega(\sqrt{dT})$ No
Lipschitz bandit $\Omega\left(T^{\frac{d+1}{d+2}}\right)$ Yes [Kleinberg et al., 2019]
Bayesian MAB ${\Omega\left(\sqrt{AT\frac{I(\pi^{*};\mathcal{H}_{T})}{\log K}}\right)}$ First bound
Bayesian contextual MAB ${\Omega\left(\sqrt{CAT\frac{I(\pi^{*};\mathcal{H}_{T})}{\log K}}\right)}$ First bound
Bayesian linear bandit $\Omega\left(\sqrt{dT\frac{I(\pi^{*};\mathcal{H}_{T})}{\log K}}\right)$ First bound
3 Regret Lower Bounds Using Fano’s Inequality
We now introduce a novel approach for obtaining worst-case Bayesian regret lower bounds, which can be applied to a wide variety of interactive decision-making tasks. This method can be viewed as an extension of Fano’s inequality for non-exact recovery [Scarlett and Cevher, 2019] to online problems. We define worst-case Bayesian regret in the following manner,
$\inf_{\{\pi_{t}\}_{t=1}^{T}}\sup_{\nu\in\Delta(\mathcal{M}\times\mathcal{C})}\mathcal{BR}_{\nu}(T;\{\pi_{t}\}_{t=1}^{T})\equiv\mathcal{BR}^{*}(T).$
To lower bound the worst-case Bayesian regret we use Fano’s inequality.
Theorem 3.1 (Fano’s inequality, Theorem 2.10 of [Cover, 1999]).
Let $X,Y\sim Q$ be two jointly distributed random variables, where $X$ can take values over a finite set, whose cardinality is $\mathcal{X}$. Let $\hat{X}=f(Y)$ for some $f$ be an estimator of $X$. Then the following holds for all $f$,
$\mathbb{P}(X\neq\hat{X})\geq 1-\frac{I(X;Y)+1}{\log\mathcal{X}}.$ (1)
Fano’s inequality can be used to lower-bound the error in hypothesis testing problems. Thus, to utilize Fano’s inequality, we need to reduce our problem from regret minimization to hypothesis testing. We assume access to some set $\Phi=\{\phi_{1},\dots,\phi_{K}\}$ of (possibly stochastic) decisions such that for some $\epsilon>0$, $\mathbb{E}_{M\sim P}[\left|\mu_{M}(\phi_{i})-\mu_{M}(\phi_{j})\right|]\geq\epsilon$, for all $i\neq j$. From Markov’s inequality, we have that for any $\phi\in\Delta(\Pi)$,
$\sup_{i}\mathbb{E}_{M\sim P}[\left|\mu_{M}(\phi_{i})-\mu_{M}(\phi)\right|]\geq\epsilon\sup_{i}\mathbb{P}\left(\left|\mu_{M}(\phi_{i})-\mu_{M}(\phi)\right|\geq\epsilon\right).$ (2)
Now, instead of lower-bounding the regret, we lower-bound the error probability. This reduces our problem from regret minimization to hypothesis testing. We also assume that a metric $\rho$ exists such that $\mathbb{E}[\left|\mu_{M}(\psi)-\mu_{M}(\phi)\right|]\geq\rho(\phi,\psi)$ for all $\psi,\phi\in\Delta(\Pi)$. Using Theorem 3.1, we can lower-bound the prediction error. This results in the following proposition.
Proposition 3.2.
For any algorithm, prior $P$, a decision set $\Phi_{K}=\{\phi_{1},\dots,\phi_{K}\}\subseteq\Delta(\Pi)$ and for all $\epsilon>0$ such that $\rho(\phi_{i},\phi_{j})\geq\epsilon$ for all $i\neq j$,
$\frac{\mathcal{BR}_{P}(T)}{T\epsilon}\geq 1-\frac{I(V;\mathcal{H}_{T})+1}{\log K}$ (3)
where $V$ is the index of the best decision in hindsight from the set $\Phi$.
Proposition 3.2 allows us to lower-bound the Bayesian regret using a specific set of decisions. We suggest a method for selecting the decision set $\Phi$, which provides a general approach for obtaining regret lower bounds. We divide our analysis into finite, parametric with infinite decisions, and non-parametric decision spaces. To make the lower bound in Equation 3 tighter, we should select the largest value of $K$ that satisfies the requirements for a given $\epsilon>0$. Thus, for finite decision spaces, we choose $\Phi$ to be an $\epsilon$-local packing set of the decision simplex, $\Delta(\Pi)$, according to $\rho$, under $P$. Thus, we have $K={\mathcal{M}}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)$. Similarly, for parametric decision spaces with infinite decisions, we choose $\Phi$ to be an $\epsilon$-local packing set of the decision set, $\Pi$, i.e., $K={\mathcal{M}}^{\mathrm{loc}}(\rho,\Pi,\epsilon)$. Yang and Barron [1999] have obtained a relationship between the local and global packing numbers, which we state in Lemma 3.3.
Lemma 3.3 (Lemma 3 of Yang and Barron [1999]).
For any space $A$ with metric $d$ and for all $\epsilon>0$,
$\log{\mathcal{M}}\left(d,A,\frac{\epsilon}{2}\right)-\log{\mathcal{M}}(d,A,\epsilon)\leq\log{\mathcal{M}}^{\mathrm{loc}}(d,A,\epsilon).$
For non-parametric decision spaces, we use an $\epsilon$-global packing set of our decision space $\Pi$, which means that $K=\mathcal{M}(\rho,\Pi,\epsilon)$. After selecting $\Phi$, we need to handle $I(V;\mathcal{H}_{T})$. This can be achieved using the following theorem.
Theorem 3.4 (Yang and Barron [1999]).
Under the assumptions and notations of Proposition 3.2, if $\Pi$ is a parametric decision space then
$I(V;\mathcal{H}_{T})\lesssim 2\epsilon^{2}T.$
If $\Pi$ is a non-parametric decision space then
$I(V;\mathcal{H}_{T})\leq\inf_{\delta>0}\left(\log\mathcal{N}_{2}\left(\Pi,\sqrt{\delta}\right)+T\delta\right).$
Theorem 3.4 allows us to obtain upper bounds for the mutual information. Furthermore, it provides us with a relationship between the global and local packing numbers. Combining Theorem 3.4, Proposition 3.2 and the selection of $\Phi$, results in the following Proposition.
Theorem 3.5.
Let there be a Bayesian interactive decision-making problem as defined in Section 2, and let $\epsilon>0$. If $\Pi$ is a finite decision space,
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\gtrsim 1-\frac{2\epsilon^{2}T+1}{\log\mathcal{M}(\rho,\Delta(\Pi),\epsilon/2)-\log\mathcal{M}(\rho,\Delta(\Pi),\epsilon)}.$ (4)
If $\Pi$ is an infinite parametric decision space,
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\gtrsim 1-\frac{2\epsilon^{2}T+1}{\log\mathcal{M}(\rho,\Pi,\epsilon/2)-\log\mathcal{M}(\rho,\Pi,\epsilon)}.$ (5)
If $\Pi$ is a non-parametric decision space,
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\geq 1-\frac{\inf_{\delta>0}\left(\log\mathcal{N}_{2}\left(\Pi,\sqrt{\delta}\right)+T\delta\right)+1}{\log\mathcal{M}(\rho,\Pi,\epsilon)}.$ (6)
To obtain regret lower bounds, all we need are the values of the packing and covering numbers of the decision space. Then, we simply need to select a value of $\epsilon$ such that the left-hand side of Equations 4, 5 or 6 is $\frac{1}{2}$ (or any other positive constant). In Appendix C we demonstrate this, by deriving regret lower bounds for several known online problems, which are summarized in Table 1. Moreover, this method can be used to obtain lower bounds for the frequentist regret, due to the mini-max theorem, which states that the worst-case Bayesian regret is equal to the mini-max frequentist regret [Lattimore and Szepesvári, 2019]. More detailed proofs and explanations are provided in Appendix C.
4 Information Theoretic Bayesian Regret Lower Bounds
Russo and Van Roy [2014, 2016] showed that the regret of Thompson sampling and IDS can be upper bounded by the mutual information.
Theorem 4.1 (Propositions 2,4 of [Russo and Van Roy, 2014] and Propositions 1,5 of [Russo and Van Roy, 2016])).
In the multi-armed bandit with $K$ arms setting, the regret of Thompson sampling and information-directed sampling can be upper bounded by $O(\sqrt{KTI(\pi^{*};\mathcal{H}_{T})})$. Furthermore, if the rewards are a linear function of the decisions, and the decisions are vectors in $\mathbb{R}^{d}$, then the regret is upper bounded by $O\left(\sqrt{dTI(\pi^{*};\mathcal{H}_{T})}\right)$.
Intuitively, in information-theoretic terms, mutual information quantifies the amount of information one random variable contains about another one [Cover, 1999]. This means that mutual information can be utilized to measure the information that is aggregated from various sources using a common measure of bits. Theorem 4.1 shows that we can upper bound the regret when provided with an upper bound for that number of bits. We can now ask the converse question: What is the lower bound of the worst-case Bayesian regret, given $I(\pi^{*};\mathcal{H}_{T})\leq R$? The following proposition answers this question, for problems with a finite decision space.
Proposition 4.2.
Let $R\leq H(\pi^{*})$ be given. Assume that the Bayesian decision problem has a finite decision space of size $K$, such that $I(\pi^{*};\mathcal{H}_{T})\leq R$. Then the worst-case Bayesian regret can be lower bounded by $\Omega\left(\sqrt{TK\frac{R}{\log K}}\right)$. If $\ \Pi\subset\mathbb{B}_{2}^{d}$ and the rewards are a linear function of the decisions, then the worst-case Bayesian regret can be lower bounded by $\Omega\left(\sqrt{Td\frac{R}{\log K}}\right)$.
The proof of Proposition 4.2 utilizes Lemma D.3. This lemma introduces an upper bound for the mutual information, similar to Theorem 3.4, that takes into account the constraint. The rest of the proof follows a similar structure to the proof of Theorem 3.5, substituting the tighter bound into Equation 4. This yields the following Proposition.
Proposition 4.3.
Let $R\leq H(\pi^{*})$ be given. Assume that the Bayesian decision problem has a finite decision space of size $K$, such that $I(\pi^{*};\mathcal{H}_{T})\leq R$. Then,
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\gtrsim 1-\frac{2\epsilon^{2}T\frac{R}{\log K}+1}{\log\mathcal{M}_{1}(\Delta(\Pi),\epsilon/2)-\log\mathcal{M}_{1}(\Delta(\Pi),\epsilon)}$ (7)
The rest of the proof follows by selecting the appropriate value of $\epsilon$. Detailed proofs are found in Appendix D. Now, we can quantify the relationship between information and regret using Theorem 4.1 and Proposition 4.2. Proposition 4.2 states that when more bits of information can be accumulated, due to the properties of the environment or the agent, more regret will be suffered. We now consider two agents in a contextual MAB environment, with $K$ arms. The first agent ignores the contextual information and plays each context independently. Thus, for this agent, the mutual information between the optimal policy $\pi^{*}$ and the history $\mathcal{H}_{T}$ is $I_{1}(\pi^{*};\mathcal{H}_{T})\leq\log K$. The second agent utilizes the information in an efficient manner, which reduces the uncertainty in the optimal arm, leading to $I_{2}(\pi^{*};\mathcal{H}_{T})\leq 1$. The amount of contextual information provided is $\log\left(\frac{K}{2}\right)$ bits. From Proposition 4.2 and Theorem 4.1 we know that the optimal behavior of the regret is $\Theta\left(\sqrt{TI(\pi^{*};\mathcal{H}_{T}})\right)$ (omitting terms that depend on the size of the decision space). Thus, the regret of the first agent is $\sqrt{T\log K}$ while for the second one, it is $\sqrt{T}$. Thus, $\log\left(\frac{K}{2}\right)$ bits of external information can be traded with $\sqrt{T}\left(\sqrt{\log K}-1\right)$ regret.
(a) Bayesian regret
(b) Accumulated information
Figure 2: (a) The Bayesian regret and (b) the accumulated information in bits for Thompson sampling under three different priors over the same bandit structure. The entropy of each prior is described in the legend. The shadowed areas correspond to 2-sigma error bars.
We can also quantify the relationship between information accumulation and regret. Suppose we have two algorithms in the MAB setting, with 16 arms and a prior such that $H(\pi^{*})=4$. During the first 10 episodes, one algorithm accumulates 2 bits of information, while the other accumulates a single bit. In the subsequent 100 episodes, both algorithms accumulate the remaining bits. Since the first algorithm accumulated 2 bits initially, it suffers a regret of $\sqrt{200}$ in these 100 episodes. The second algorithm suffers a regret of $\sqrt{300}$ since it needs to accumulate 3 additional bits. Thus, the second algorithm suffers more regret, even though both accumulate the same number of bits by the end.
5 Experiments
5.1 Stochastic Bayesian bandit
We begin by demonstrating the trade-off in a simple Bayesian MAB problem. In this problem, our model space contains $K$ possible models. The decision space of the problem also contains $K$ decisions, which are called arms. The mean of all arms in every model is $0.5(1-\varepsilon)$, except for one arm for which the mean is $0.5(1+\varepsilon)$. The optimal arm is different for every model. The reward of each arm is Bernoulli distributed. We also fix $K=8,\varepsilon=0.1$. We choose this setting since all of the problems have the same structure. Thus, any difference in regret stems from differences in the prior. In the first experiment, we compare the Bayesian regret obtained by Thompson sampling with different priors for every problem. This experiment allows us to isolate the effect of the total bits on the regret. Results are presented in Figure 2. We see that when fewer bits of information can be accumulated, fewer regret is suffered.
The next experiment is performed under the same setting. Now we assume a uniform prior and compare three different bandit algorithms - EXP3 [Lattimore and Szepesvári, 2020], APS [Xu and Zeevi, 2023], and Thompson sampling [Thompson, 1933]. We also estimate the bits of accumulated information for each algorithm. Results are presented in Figure 3. We can now see how measuring the accumulated bits of information also quantifies the relationship in regret between different bandit algorithms. Additional experimental details are provided in Appendix E.
(a) Bayesian regret
(b) Accumulated information
Figure 3: (a) The Bayesian regret and (b) the accumulated information in bits for three different bandit algorithms, under the same bandit structure and a uniform prior. The bandit algorithms used are described in the legend. The shadowed areas correspond to 2-sigma error bars.
5.2 Question Answering with Large Language Models
In the following experiment, we show how the regret-information trade-off can be utilized in a practical setting. We consider a sequential multiple choice question-answering (MCQA) task, where every round we need to provide an answer for a question, out of four possible answers. We also have access to two large language models (LLM), where one has significantly more parameters than the other. Every turn we need to choose which LLM should answer the question. As contextual information, every round we are provided with the question, possible answers, and the output of the small LLM for the given prompt. We receive a positive reward of 1 for answering the question correctly. If we choose to deploy the large LLM we incur a negative reward of 0.1. Our goal is to minimize the accumulated regret.
Since the LLM outputs a probability distribution over tokens, we can measure the information the small LLM will gain after answering the question, in bits. We prompt the small LLM with the question and possible answers, which returns scores for each token. We take the scores corresponding only for the tokens [A, B, C, D], each corresponding to a different answer, and use them to get a distribution over these tokens. The number of information bits is the KL divergence between this distribution and the uniform one. Now we create a bits-based policy, which deploys the large LLM only if the mutual information is below some threshold. This policy queries the large LLM only when a substantial amount of bits can be acquired. We compare the mean regret over a horizon of 200 steps between this policy, and one that randomly selects which LLM to use. The selected threshold value for the bits-based policy also ensures that we query the large LLM the same number of times as the random policy.
We run the experiment described above with the following specifications. To generate the multiple-choice questions, we used the MMLU dataset, which contains multiple-choice questions in a variety of topics, such as algebra, business, etc. [Hendrycks et al., 2021a, b].
Table 2: Mean regret for both policies for a horizon of 200 episodes, under different deployment percentages of the large LLM. The left column denotes the percentage of queries to the large LLM. We report the standard error of the mean.
% Large LLM Bits-based Random
$50\%$ $\mathbf{30.1\pm 0.5}$ $32.5\pm 0.2$
$70\%$ $\mathbf{22.7\pm 0.5}$ $24.8\pm 0.1$
$90\%$ $\mathbf{15.5\pm 0.3}$ $17.1\pm 0.2$
Figure 4: Regret of random and bits-based selection in the MCQA with LLMs task. The random and bits-based policies were allowed to use the large LLM for only 90% of the episodes. Both policies were evaluated across 10 different seeds, with the same questions. The shadowed areas correspond to 2-sigma error bars.
200 questions were randomly selected in advance, which we used to evaluate both agents with 10 different seeds while maintaining the same questions and order. We then repeated this process for 10 different randomly generated sets of questions. We used Mistral 7B Jiang et al. [2023] for the small LLM, and Mixtral 7Bx8 [Jiang et al., 2024] for the large one. For both models, 4-bit quantization [Dettmers and Zettlemoyer, 2023] and flash attention Dao et al. [2022] were applied. Figure 4 compares the random and the bits-based policy for one randomly generated set of questions. Table 2 describes the mean regret of every policy under a different number of large LLM deployments. From the results, we see that deciding whether to query the large LLM or not using bits is better than random selection. Furthermore, this improvement becomes more significant as we increase the number of deployments allowed. This demonstrates how we can easily utilize the quantification of information in bits in online settings. Additional details are provided in Appendix E. Code is available at https://github.com/itaishufaro/bitsandbandits.
6 Related Work
Bayesian setting:
In the Bayesian setting, prior information is assumed to be expressed as a probability distribution, called the prior. New information is then incorporated using Bayesian inference to update the prior probability, resulting in a posterior distribution. Bayesian algorithms have been extensively studied in multiple online decision-making problems such as multi-armed bandits [Russo and Van Roy, 2014, Kaufmann et al., 2012] and reinforcement learning [Guez et al., 2012, Ghavamzadeh et al., 2016]. Upper bounds for Bayesian algorithms have been proved for both the frequentist [Kaufmann et al., 2011] and Bayesian settings [Russo and Van Roy, 2014, 2016]. Prior dependent regret bounds for Bayesian algorithms have also been studied [Bubeck and Liu, 2013, Russo and Van Roy, 2014]. Previous studies have explored prior-dependent lower bounds for Bayesian regret, but these are constrained to specific forms of priors, such as Gaussian priors [Atsidakou et al., 2024].
Contextual information:
In the contextual setting, information is revealed to the agent before every round, which can be leveraged to minimize regret. This framework has gathered attention within both the bandit [Bietti et al., 2021] and reinforcement learning frameworks [Klink et al., 2020, Modi and Tewari, 2020]. The contexts can be selected arbitrarily by an adversary [Beygelzimer et al., 2011, Chu et al., 2011] or generated from some prior probability distribution, similarly to the Bayesian setting [Hao et al., 2020, May et al., 2012]. Bayesian algorithms have also been adapted to the arbitrary context setting [Agrawal and Goyal, 2013]. Furthermore, the type of information provided to the learner can expand beyond the contextual information provided directly [Schneider and Zimmert, 2024]. Our work introduces a contextual Bayesian setting framework that covers a wide variety of interactive decision-making tasks.
Information-theoretic methods
have been employed to derive upper bounds for various online tasks. Russo and Van Roy [2016] introduced an information-theoretic method for establishing Bayesian regret upper bound for Thompson sampling in the multi-armed bandit setting, which depends on the entropy of the prior. This approach was also used to design new Bayesian algorithms that utilize mutual information called information-directed sampling (IDS) [Russo and Van Roy, 2014]. IDS and the information-theoretic method it utilizes were also extended to other tasks such as contextual bandits [Neu et al., 2022], sparse linear bandits [Hao et al., 2021], non-stationary bandits [Min and Russo, 2023, Liu et al., 2023], reinforcement learning [Lu and Van Roy, 2019], non-linear control [Kakade et al., 2020], partial monitoring [Lattimore and Szepesvári, 2019], and other online optimization problems [Liu et al., 2018, Dong et al., 2019, Lattimore and Gyorgy, 2021, Russo and Van Roy, 2018]. The bounds presented in works such as [Russo and Van Roy, 2016, 2014, Neu et al., 2022] are based on the upper bounding of the information ratio. Our work uses a different approach, which obtains nearly matching prior-dependent lower bounds for Thompson sampling and information-directed sampling (IDS). Furthermore, these bounds can be applied beyond the bandit setting. Seldin et al. [2011] used a different information-theoretic method. They utilized PAC-Bayes bounds to obtain information-theoretic upper bounds on the per-round regret, that depend on mutual information.
7 Conclusion and Future Work
We introduce a general setting that embeds contextual information in a Bayesian setting. We showed that the regret can be lower bounded by $\Omega\left(\sqrt{TKI(\pi^{*};\mathcal{H}_{T})/\log K}\right)$. Combining this result with known upper bounds, this quantifies the relationship between a-priori external information and regret. Furthermore, in online settings these bounds quantify the relationship between online information accumulation and regret. We also utilized this trade-off in an MCQA task with LLMs, demonstrating that the aforementioned quantification can be easily used in an online setting.
The general setting we introduced is relevant not only to the bandit setting, but also to general sequential decision problems. A central assumption in our analysis is that exogenous information can be gathered regardless of the process history. However, this assumption may be violated in general sequential decision process where the action itself may have ramifications on the quality, quantity and cost of the exogenous information.
Our work only covers the Bayesian setting, and our definition of the measure of information relies on it. In other settings, such as adversarial learning [Neu and Olkhovskaya, 2020] or frequentist settings. Finally, and importantly, with the increasing prevalence of LLMs, and foundation models in general, building solid foundations as well as practical algorithms for using prior knowledge in sequential decision-making is an important research endeavor that may be built upon the foundations that are laid in this paper.
Acknowledgements
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101034255.
Appendix A Table of Notations
Table 3: Commonly used notations throughout the paper and the appendix.
Notations Description
$\Pi$ The decision space.
$\mathcal{M}$ The model class.
$\mathcal{O}$ The observation space.
$\mathcal{C}$ The context space.
$\pi$ A decision.
$R_{M}(\pi)$ reward distribution according to decision $\pi$ on model $M$.
$\mu_{M}(\pi)$ mean reward according to decision $\pi$ on model $M$.
$P$ Prior probability of the task.
$\pi^{*}_{M}$ The optimal decision for a model $M$.
$\mathcal{H}_{t}$ The total accumulated history up to round $t$
(including observations and contexts).
$p_{t}$ Policy selected at round $t$.
$r_{t}$ Reward sampled at time $t$.
$\rho$ A norm that lower bounds the difference in reward means.
Appendix B Useful Properties of Covering Numbers
Definition B.1 (Covering and packing numbers).
Let $H$ be a normed space, and let $d(\cdot,\cdot)$ be the metric induced by the norm. Let $\epsilon>0$.
  • A set $C\subset H$ is said to be a covering set of $H$ if for all $h\in H$ there is $c\in C$ such that $d(h,c)\leq\epsilon$.
  • The covering number is defined as
    $\mathcal{M}(d,\epsilon,H)=\max\left\{m:\exists C,\left|C\right|=m\text{ and $C$ is a covering set of $H$}\right\}.$
    If $d(x,y)=\left\|x-y\right\|_{p}$, we denote it by $\mathcal{M}_{p}(\epsilon,H)$.
  • A set $C\subset H$ is said to be a packing set of $H$ if for all $h\in H$ there is $c\in C$ such that $d(h,c)\geq\epsilon$.
  • The packing number is defined as
    $\mathcal{N}(d,\epsilon,H)=\min\left\{m:\exists C,\left|C\right|=m\text{ and $C$ is a packing set of $H$}\right\}.$
    If $d(x,y)=\left\|x-y\right\|_{p}$, we denote it by $\mathcal{N}_{p}(\epsilon,H)$.
  • We now define $B(\theta,\epsilon)=\{\theta^{\prime}\in H:d(\theta^{\prime},\theta)\leq\varepsilon\}$. $C$ is an $\epsilon$ local packing set of $B(\theta,\epsilon)$ if $\forall\theta^{\prime}\in B$, $\exists\theta^{\prime\prime}\in C$ such that $d(\theta^{\prime},\theta^{\prime\prime})\leq\frac{\epsilon}{2}$.
  • The local packing number is defined as
    $\mathcal{M}^{\mathrm{loc}}(d,\epsilon,H)=\max\left\{m:\exists B(\theta,\varepsilon),\text{such that $C$ is an }\epsilon-\text{covering of $B$ and }\left|C\right|=m\right\}.$
Lemma B.1.
The following properties hold:
  • (Theorem 14.2 of Wu [2020]) For any subset of dimension $d$ that is contained in $\mathbb{B}_{1}^{d}$ the covering number $\mathcal{N}_{2}(\epsilon)$ obeys $\log\mathcal{N}_{2}(\epsilon)=\Theta(d\log(1/\epsilon))$. This includes the probability simplex with $d$ variables and $\mathbb{B}_{d}^{2}$.
  • (Lemma 6 of Gottlieb et al. [2017]) The covering of functions from $[0,1]^{d}$ to $[0,1]$ that are $L$-Lipschitz is given by $\log\mathcal{N}_{2}(\epsilon)=\Theta\left((L/\epsilon)^{d}\right)$.
  • (Theorem 14.1 of Wu [2020]) Under the same norms, for the same set,
    $\mathcal{M}(\epsilon/2)\leq\mathcal{N}(\epsilon)\leq\mathcal{M}(\epsilon)$
Appendix C Proofs for Section 3
C.1 Proof of Proposition 3.2
Proof.
We write
$\mathcal{BR}^{*}(T)$ $=\sup_{\nu}\inf_{\pi_{1},\dots,\pi_{T}}\mathbb{E}_{M\sim\nu}\left[\sum_{t=1}^{T}(\mu_{M}(\pi^{*})-\mu_{M}(\pi_{t}))\right]$
$\geq\sup_{\nu}\inf_{\pi_{1},\dots,\pi_{T}}\epsilon T\mathbb{P}_{\nu}\left(\sum_{t=1}^{T}(\mu_{M}(\pi^{*})-\mu_{M}(\pi_{t}))\geq\epsilon T\right)$
$\geq\epsilon T\sup_{\nu}\inf_{\pi_{1},\dots,\pi_{T}}\mathbb{P}_{\nu}\left(\forall t:\mu_{M}(\pi^{*})-\mu_{M}(\pi_{t})\geq\epsilon\right)$
$\geq\epsilon T\sup_{\nu}\sup_{k\in\{1,\dots,K\}}\mathbb{P}\left(\forall t:\left|\mu_{M}(\phi_{k})-\mu_{M}(\pi_{t})\right|\geq\epsilon\right)$
$\geq\epsilon T\sup_{\nu}\sup_{k\in\{1,\dots,K\}}\mathbb{P}\left(\forall t:\rho(\phi_{k},\pi_{t})\geq\epsilon\right)$
$\geq\epsilon T\left(1-\frac{I(V;\mathcal{H}_{T})+1}{\log K}\right)\,,$
where the last inequality follows from Fano’s inequality (Theorem 3.1). ∎
C.2 Proof of Theorem 3.5
Proof.
We provide proof for three different scenarios.
  • First scenario, finite $\Pi$. For this scenario, we choose $\Phi$ to be a local packing set of the decisions simplex $\Delta(\Pi)$, so $K=\mathcal{M}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)$. Applying this selection of set, and using Equation 3 we have that
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq 1-\frac{1+I(V;\mathcal{H}_{T})}{\log\mathcal{M}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)}.$ (8)
    Applying Theorem 3.4 we have that $I(V;\mathcal{H}_{T})\leq 2\epsilon^{2}T$. Furthermore, from Lemma 3.3, we have that $\log\mathcal{M}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)\geq\log\mathcal{M}(\rho,\Delta(\Pi),\epsilon/2)-\log\mathcal{M}(\rho,\Delta(\Pi),\epsilon)$. Substituting this back into Equation 8,
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq 1-\frac{1+2\epsilon^{2}T}{\log\mathcal{M}(\rho,\Delta(\Pi),\epsilon/2)-\log\mathcal{M}(\rho,\Delta(\Pi),\epsilon)}.$
  • Second scenario, infinite-parametric $\Pi$. For this scenario, we choose $\Phi$ to be a local packing set of the decision space $\Pi$, so $K=\mathcal{M}^{\mathrm{loc}}(\rho,\Pi,\epsilon)$. Applying this selection of set, and using Equation 3 we have that
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq 1-\frac{1+I(V;\mathcal{H}_{T})}{\log\mathcal{M}^{\mathrm{loc}}(\rho,\Pi,\epsilon)}.$ (9)
    Applying Theorem 3.4 we have that $I(V;\mathcal{H}_{T})\leq 2\epsilon^{2}T$. Furthermore, from Lemma 3.3, we have that $\log\mathcal{M}^{\mathrm{loc}}(\rho,\Pi,\epsilon)\geq\log\mathcal{M}(\rho,\Pi,\epsilon/2)-\log\mathcal{M}(\rho,\Pi,\epsilon)$. Substituting this back into Equation 9,
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq 1-\frac{1+2\epsilon^{2}T}{\log\mathcal{M}(\rho,\Pi,\epsilon/2)-\log\mathcal{M}(\rho,\Pi,\epsilon)}.$
  • Third scenario, non-parametric $\Pi$. For this scenario, we choose $\Phi$ to be a global packing set of the decision space $\Pi$, so $K=\mathcal{M}(\rho,\Pi,\epsilon)$. Applying this selection of set, and using Equation 3 we have that
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq 1-\frac{1+I(V;\mathcal{H}_{T})}{\log\mathcal{M}(\rho,\Pi,\epsilon)}.$ (10)
    Applying Theorem 3.4 we have that $I(V;\mathcal{H}_{T})\leq\inf_{\delta>0}\left(\log\mathcal{N}_{2}\left(\Pi,\sqrt{\delta}\right)+T\delta\right)$. Substituting this back into Equation 10,
    $\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\geq 1-\frac{1+\inf_{\delta>0}\left(\log\mathcal{N}_{2}\left(\Pi,\sqrt{\delta}\right)+T\delta\right)}{\log\mathcal{M}(\rho,\Pi,\epsilon)}.$
C.3 Proofs for Additional Lower Bounds
Lemma C.1.
For any algorithm, there exists a Bayesian decision-making task with a finite decision such that the regret can be lower bounded by $\Omega(\sqrt{\left|\Pi\right|T})$.
Proof.
We focus on the set $\Delta(\Pi)$ and note that it is a parametric set. We also see that for all decisions $\phi,\psi\in\Pi$ we have that
$\mathbb{E}[\left|\mu_{M}(\phi)-\mu_{M}(\psi)\right|]\gtrsim\left\|\phi-\psi\right\|_{1}.$
Hence, in this scenario, we can take $\rho(x,y)=\left\|x-y\right\|_{1}$. Furthermore, using Lemma B.1 we have that
$\log\mathcal{M}_{1}\left(\epsilon/2,\Delta(\Pi)\right)-\log\mathcal{M}_{1}\left(\epsilon,\Delta(\Pi)\right)=\Theta\left(\left|\Pi\right|\right).$
Substituting this into Equation 4 we obtain that
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\gtrsim 1-\frac{1+2\epsilon^{2}T}{\left|\Pi\right|}.$
Now we fix $\epsilon$ such that $\frac{2\epsilon^{2}T}{\left|\Pi\right|}=\frac{1}{2}$. More specifically, we set $\epsilon=\frac{1}{4}\sqrt{\frac{\left|\Pi\right|}{T}}$ which yields
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\gtrsim\frac{1}{2}\sqrt{\left|\Pi\right|T}$
which concludes our proof. ∎
The results for multi-armed bandits, contextual bandits, and tabular MDPs are shown by substituting $\left|\Pi\right|=A,CA$ and $SAH$ respectively.
Lemma C.2.
For any algorithm, the exists a Bayesian decision-making task with $\Pi\subseteq\mathbb{B}_{2}^{d}$, where the reward means are a linear function of the decisions, such that the regret can be lower bounded by $\Omega(\sqrt{dT})$.
Proof.
Similarly to Lemma C.1, we have $\rho(x,y)=\left\|x-y\right\|_{1}$. We focus on the set $\mathbb{B}_{2}^{d}$ and note that it is a parametric set. Furthermore, using Lemma B.1 we have that
$\log\mathcal{M}_{1}\left(\epsilon/2,\mathbb{B}_{2}^{d}\right)-\log\mathcal{M}_{1}\left(\epsilon,\mathbb{B}_{2}^{d}\right)=\Theta\left(d\right).$
Substituting this into Equation 5 we obtain that
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\gtrsim 1-\frac{1+2\epsilon^{2}T}{d}.$
Now we fix $\epsilon$ such that $\frac{2\epsilon^{2}T}{d}=\frac{1}{2}$. More specifically, we set $\epsilon=\frac{1}{4}\sqrt{\frac{d}{T}}$ which yields
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\gtrsim\frac{1}{2}\sqrt{dT}$
which concludes our proof. ∎
In the Lipshitz bandit setting, $\Pi$ is some metric space with metric $\rho$ and $\mathcal{M}=\mathcal{M}_{\mathcal{F}}$, where
=$\mathcal{F}=\{f:\Pi\to[0,1]|f\text{ is $1$-Lipschitz w.r.t }\rho\}.$
Furthermore, we assume that there exists some $d$ such that $\mathcal{M}_{\rho}(\epsilon,\Pi)\gtrsim\epsilon^{-d}$. Additionally, we know from Lemma B.1 that $\log\mathcal{N}_{2}(\epsilon,\Pi)\lesssim\epsilon^{d}$. We now provide a regret lower bound for Lipschitz bandits.
Lemma C.3.
The regret lower bound for the Lipschitz bandit is $\Omega\left(T^{\frac{d+1}{d+2}}\right)$.
Proof.
Due to Lemma B.1, the covering and packing number can now be bounded by
$\log\mathcal{N}_{2}\left(\sqrt{\frac{2}{\delta}},\Pi\right)$ $\lesssim\left(\frac{2}{\delta}\right)^{d/2}$
$\log\mathcal{M}_{\rho}(\epsilon,\Pi)$ $\gtrsim{\epsilon}^{-d}.$
Substituting this into Equation 6,
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}$ $\gtrsim 1-\left({1+\left(\frac{2}{\delta}\right)^{d/2}+T\delta}\right)\epsilon^{d}.$
We choose $\delta=2^{\frac{d}{d+2}}T^{-\frac{2}{d+2}}$. This yields that
$\frac{\mathcal{BR}^{*}(T)}{\epsilon T}\gtrsim 1-\left(1+(2T)^{\frac{d}{d+2}}\right)\epsilon^{d}.$
Selecting $\epsilon=\left(2T\right)^{-\frac{1}{d+2}}4^{-\frac{1}{d}}$ yields
$\mathcal{BR}^{*}(T)$ $\gtrsim 2^{-\frac{1}{d+2}}4^{-\frac{1}{d}}T^{\frac{d+1}{d+2}}-o(1)$
which yields the expected (and optimal) bound. ∎
We now present the following theorem, which shows that the worst-case Bayesian regret is equal to the mini-max frequentist regret.
Theorem C.4 (Theorem 1 of Lattimore and Szepesvári [2019]).
$\inf_{\{\pi_{i}\}_{i=1}^{T}}\sup_{r_{1},\dots,r_{t}}\max_{\pi^{*}\in\Pi}\mathbb{E}\left[\sum_{t=1}^{T}\left(r_{t}(\pi^{*})-r_{t}(\pi_{t})\right)\right]=\sup_{\nu}\inf_{\{\pi_{i}\}_{i=1}^{T}}\max_{\pi^{*}\in\Pi}\mathbb{E}\left[\sum_{t=1}^{T}\left(r_{t}(\pi^{*})-r_{t}(\pi_{t})\right)\right]$
This theorem then states that, without any constraints on the prior, the worst-case Bayesian regret is equal to the mini-max regret. This means that any lower bound derived for the Bayesian regret can also be applied to the frequentist one.
Appendix D Proofs for Section 4
We begin by defining the W-Lambert function and supplying useful properties of it.
Definition D.1.
The Lambert W-function is $W:[-1,\infty]\to[-1,\infty]$ is the inverse function to $x2^{x}$. This means that for all $x\geq 0$, $y=x2^{x}$ if and only if $W(y)=x$.
We now state the following useful property of the W-Lambert function.
Lemma D.1 (Equation 21 of [Iacono and Boyd, 2017]).
$W(x)=\Theta\left(\log\left(\frac{x}{\log x}\right)\right)$
We begin by proving the following useful lemma.
Lemma D.2.
Assume a finite decision space with $n$ decisions. Then the highest $L_{1}$ norm of every decision in the set, under the mutual information constraint of $I(\pi^{*};\mathcal{H}_{T})\leq R\leq H(\pi^{*})$ is at most $\Theta\left(\frac{R}{\log n}\right)$.
Proof.
We begin by noting that the mutual information can be written as $I(\pi^{*};\mathcal{H}_{T})=D\left(P^{\pi^{*}\mid\mathcal{H}_{T}},P^{\pi^{*}}\right)$. Thus, the constrained decision space becomes the decisions space under the KL constraint, which can be written as
$\sum_{i}\phi(i)\log\frac{\phi(i)}{P(i)}\leq R.$
Thus, we have that
$\phi^{*}\log\frac{\phi^{*}}{p}\leq R$
we re-parameterize $\psi=\frac{\phi^{*}}{p}$ and obtain
$p\psi\log\psi\leq R.$
Hence,
$\psi\leq 2^{W(R/p)}=\Theta\left(\frac{R}{p\log\left(R/p\right)}\right).$
We substitute $\phi^{*}$ into the expression,
$\phi^{*}\leq\Theta\left(\frac{R}{\log\left(R/p\right)}\right)=\Theta\left(\frac{R}{\log n}\right).$
We now state the following lemma that explains how we can use the results of Lemma D.2 to Equation 4.
Lemma D.3.
Under the assumptions and notations of Proposition 3.2, and under the constraint of $I(\pi^{*};\mathcal{H}_{T})\leq R\leq H(\pi^{*})$, the following holds.
$I(V;\mathcal{H}_{T})\lesssim 2T\frac{R}{\log K}\epsilon^{2}.$
Proof.
The proof follows a similar analysis to the proof of Theorem 3.4, which is provided in Yang and Barron [1999]. We denote $E$ as the space of the constrained distributions. We now have
$I(V;\mathcal{H}_{T})$ $\leq T\max_{p_{1},p_{2}\in E}D\left(p_{1},p_{2}\right)$
$\lesssim T\frac{R}{\log K}\max_{p_{1},p_{2}\in E}\left\|p_{1}-p_{2}\right\|_{2}^{2}$
$\leq T\frac{R}{\log K}\epsilon^{2}.$
where the second inequality follows from D.2. ∎
As a corollary from Lemma D.3, we obtain the following.
Corollary D.4 (Proposition 4.3).
Let there be a Bayesian interactive decision-making problem as defined in Section 2, with a finite decision space $\Pi$, where $\left|\Pi\right|=K$. Then,
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\gtrsim 1-\frac{2\epsilon^{2}T\frac{R}{\log K}+1}{\log\mathcal{M}_{1}(\Delta(\Pi),\epsilon/2)-\log\mathcal{M}_{1}(\Delta(\Pi),\epsilon)}$
Proof.
The proof follows immediately by utilizing Lemma D.3 and using the fact that in this scenario, $\rho(x,y)=\left\|x-y\right\|_{1}$, similarly to the proof in Lemma C.1. ∎
D.1 Proof of Proposition 4.2
Proof.
We begin by denoting the polytope space given the mutual information constraint as $\Delta_{R}$. We first see that
$\log\mathcal{M}_{1}(\epsilon/2,\Delta_{R})-\log\mathcal{M}_{1}(\epsilon,\Delta_{R})=\Theta\left(K\right).$
Substituting this into Equation 7,
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\gtrsim 1-\frac{2T\frac{R}{\log K}\epsilon^{2}+1}{K}.$
Selecting $\epsilon=\frac{1}{4}\sqrt{\frac{RK}{T\log K}}$ we now have that
$\mathcal{BR}^{*}(T)\gtrsim\frac{1}{2}\sqrt{TK\frac{R}{\log K}}$
We present the proof for the linear case. Now,
$\log\mathcal{M}_{1}(\epsilon/2,\Delta_{R})-\log\mathcal{M}(\epsilon,\Delta_{R})=\Theta\left(d\right).$
So now we have
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\gtrsim 1-\frac{2T\frac{R}{\log K}\epsilon^{2}+1}{d}$
selecting $\epsilon=\frac{1}{4}\sqrt{\frac{Rd}{T\log K}}$ we now have that
$\mathcal{BR}^{*}(T)\gtrsim\frac{1}{2}\sqrt{\frac{dTR}{\log K}}.$
Corollary D.5.
For a $K$-Bayesian MAB, the regret can be lower bounded by $\Omega\left(\sqrt{KT\frac{I(\pi^{*};\mathcal{H}_{T})}{\log K}}\right)$. If the MAB is also linear, then the lower bound becomes $\Omega\left(\sqrt{dT\frac{I(\pi^{*};\mathcal{H}_{T})}{\log K}}\right)$. For a Contextual MAB with $C$ contexts, the lower bound becomes $\Omega\left(\sqrt{KCT\frac{I(\pi^{*};\mathcal{H}_{T})}{\log K}}\right)$.
Appendix E Experiments Details
Figure 5: The prompting done for the LLMs in our experiments. The question is preceded by "Answer the following question:". Following this, the options are presented and to improve accuracy, the LLM is requested to output only the letter for the correct answer. Then, for the next round’s prompt, we add the LLM’s answer and feedback.
E.1 Bayesian Bandit Experiments
The mutual information for the experiments in this setting is estimated using the KL divergence between the output probability distribution of every algorithm with the prior probability, $D\left(p_{t},P^{\pi^{*}}\right)$. This is an estimator of the mutual information, similar to the one used by Seldin et al. [2011]. The mutual information was then calculated according to $D\left(p_{t},P^{\pi^{*}}\right)$, where $p_{t}$ is the real posterior. In our problem, the posterior update is done in the following manner.
$P(\pi^{*}=i\mid\mathcal{H}_{t})$ $=\frac{P(\mathcal{H}_{t}\mid i)P(i)}{P(\mathcal{H}_{t})}$
$=\frac{P(\mathcal{H}_{t}\mid i)P(i)}{\sum_{j}P(\mathcal{H}_{t}\mid i)P(i)}$
where
$P(\mathcal{H}_{t}\mid i)$ $=\left(\frac{1+\epsilon}{2}\right)^{N_{s,i}}\left(\frac{1-\epsilon}{2}\right)^{N_{f,i}}\prod_{j\neq i}\left[\left(\frac{1-\epsilon}{2}\right)^{N_{s,j}}\left(\frac{1+\epsilon}{2}\right)^{N_{f,j}}\right],$
and $N_{s_{i}}$ is the number of successful pulls (pulls that received reward 1) of arm $i$. $N_{f,i}$ is the number of failed pulls of arm $i$.
In both bandit experiments, we used 200 random seeds. Each random seed considers both the randomness in the problem generated (since the bandit problem is generated via a prior) and the randomness during the algorithm’s execution. We then reported in each graph the standard error for both the regret and mutual information, which is given by $\mathrm{err}=\frac{2\sigma}{\sqrt{n}}$ where $\sigma$ is the standard deviation and $n$ is the number of seeds.
E.2 MCQA with LLM Experiment
For the second experiment, we used two different LLMs. The first is Mistral 7B Jiang et al. [2023], an LLM with 7.3B parameters. The second is Mixtral 8x7B Jiang et al. [2024]. The architecture of this model is the same as Mistral 7B, but it uses 8 of them and a mixture of experts approach to select which one to use. For both models, we use 4bit quantization of type nf4 Dettmers and Zettlemoyer [2023]. In every round, we queried the LLM with all previous prompts along with a prompt that consists of the following structure: "Answer the following question: [question]? These are the possible answers [options]. Please output only the letter for the correct answer." Using the LLM’s output we obtain a probability distribution over the tokens of the possible answers [A,B,C,D]. We then selected the answer according to this distribution and received feedback as to whether the answer was correct or not. The feedback was also added to the prompt. The entire procedure is described in Figure 5.
We repeated this process for 10 different question seeds. For each question seed, a different set of 200 questions was selected. Both policies were evaluated on each set for 10 different seeds, so 100 random seeds were used for this experiment. Figure 4 describes the behavior of both policies across 10 different seeds but using the same question seed. Table 2 reports the results across all 100 seeds.
E.3 Compute Specifications
The bandits experiment is performed on a CPU and does not require special compute workers. The experiments using Mistral 7B were evaluated using a machine with RTX4090. The experiments using Mixtral 8x7B were done using a machine with 3xA40.
E.4 Dataset and Models License
The MMLU dataset Hendrycks et al. [2021a, b] is available under the MIT license. Mistral 7B Jiang et al. [2023] and Mixtral 8x7B Jiang et al. [2024] are both available under the Apache license 2.0.